My daughter is having a problem solving this question. I cannot figure out how to work out the problem. Please help!
a-4/5=12
Answer is 64

Answer by Alan3354(69443)

a-4/5=12
(a-4)/5 = 12 is clearer
Multiply by 5
a-4 = 60
Add 4
a = 64